MK is a 21 year old White female college student who has been bothered by a cold for most of the semester. She has several colds a year but they usually do not last as long as this one. She presented with the usual symptoms: a sore throat, blocked sinuses, and a hacking cough with moderate quantities of slightly yellowish sputum. Two days ago, her cough was considerably worse. The sputum increase was purulent. MK also had a temperature of 99.8⁰ and C/O (complained of) pain in the left side of her chest while breathing. Her respirations were shallow. She felt weak and ached all over. She felt so bad that she did not feel like eating anything. Yesterday she was getting ready to take a final exam when she collapsed in her room. Her roommate and some friends brought her to the campus infirmary. The MD listened to her breath sounds and diagnosed her with pneumonia + exhaustion. She was admitted to a local hospital. Her admission orders included SBR (strict bed rest), appropriate antibiotics, respiratory treatments and a regular diet.

Normally dietitians do not visit patients with an order for a regular diet unless they are requested to do so. However, this hospital has an excellent nutritional screening process that requires a diet technician to review admission criteria, height, weight, and admission lab values. All of the above information goes into the hospital computer on admission. The technician accomplishes the screening by simply pulling up the new admissions for the previous day on the computer and reviewing for the four parameters previously mentioned. This can be done while sitting in the dietary offices. The techs have a form to fill out as they review the computerized chart. The form will vary from hospital to hospital, but there will be a thread of commonality in all screening forms. If the form requires extensive information, not all of the information required may be obtainable from the computer. There are times when the tech will have to visit the patient to obtain additional information. The day after admission, MK was screened using the computer and a brief interview by the diet tech. The following information was obtained:

MK’s MH (medical history) was unremarkable except for frequent colds. Her weight history was very significant. She claimed that her UBW was 115 lb. and she had lost 17 lb. over the last three months. She attributed this weight loss to a very strenuous semester. The computer indicated her height and weight to be:

  • Height 5’4”
  • Weight 98 lbs.

QUESTIONS:

  1. Determine MK’s IBW
  2. Calculate her percent IBW and percent UBW. Please show work for all calculations.
  3. Calculate MK’s BMI and explain what the results mean.
Answer #1

1.MK'S IDEAL BODY WEIGHT IS 59kg

Calculation - height - 105

= height in cm - 105

=164 - 105

= 59 kg

In pounds 130 lb is the ideal body weight for MK2.

2. % ideal body weight is

Formula

Weight × 100 ÷ ibw

= 98×100÷130

9800÷130

=75.38 %IBW

% UBW ,

Formula is

100 × actual weight ÷ usual weight

=100 × 98÷115

=9800÷115

= 85.21 is % UBW for MK

3.MKs BMI

Formula

BMI = wt  in lb÷ (ht in inches × ht in inches)×703

= 98÷ (64×2)×703

16.8 it indicated under weight.